### v1

# Probabilistic Sampling of Balanced K-Means using Adiabatic Quantum Computing

Abstract
Adiabatic quantum computing (AQC) is a promising quantum computing approach for discrete and often NP-hard optimization problems. Current AQCs allow to implement problems of research interest, which has sparked the development of quantum representations for many computer vision tasks. Despite requiring multiple measurements from the noisy AQC, current approaches only utilize the best measurement, discarding information contained in the remaining ones. In this work, we explore the potential of using this information for probabilistic balanced k-means clustering. Instead of discarding non-optimal solutions, we propose to use them to compute calibrated posterior probabilities with little additional compute cost. This allows us to identify ambiguous solutions and data points, which we demonstrate on a D-Wave AQC on synthetic and real data.
Figure 1: The proposed approach uses an adiabatic quantum computer to sample solutions of a balanced k-means problem. By using an energy-based formulation, likely solutions are drawn from a Boltzmann distribution. By reparametrizing the distribution, the calibrated posterior probability of each solution can be estimated.
1 Introduction
Clustering is a fundamental problem in machine learning and computer vision, extensively employed for the analysis and organization of large volumes of unlabeled visual data. It involves grouping similar objects based on their features, facilitating efficient data processing and unsupervised learning algorithms (Xu & Wunsch, 2005). The information extracted from clusters plays a vital role in various computer vision applications, including image classification (Lazebnik et al., 2006; Caron et al., 2018), segmentation (Coleman & Andrews, 1979; Dhanachandra et al., 2015), tracking (Keuper et al., 2020), and network training (Coates & Ng, 2012; Yang et al., 2016). However, the often NP-hard computational complexity of solving clustering problems often hinders their application to large-scale problems that require fast processing times (Xu & Wunsch, 2005).
In addition to this, the ambiguity of the task itself contributes to the absence of a unique definition for clustering (Xu & Wunsch, 2005), which arises from the range of possible optimization objectives. An example of this are different algorithms that can optimize for the most compact solution, an underlying data-generating process or just use local distances. One way to address this challenge is the use of confidence-based estimators and the sampling of multiple likely solutions. This approach helps in identifying well-assigned samples and uncovering ambiguities within the data (Xu & Wunsch, 2005).
Quantum algorithms that promise a considerable speedup over their classical counterparts and are inherently probabilistic have the potential to enable a new family of machine learning algorithms. By now, quantum machine learning algorithms approach tasks such as optimization of quadratic problems (Kadowaki & Nishimori, 1998), training of restricted Boltzmann machines (Dixit et al., 2021), and learning with quantum neural networks (Abbas et al., 2021). While the current quantum computers can only solve small-scale problems, they provide the basis to develop and test algorithms that can considerably increase the size of feasible problems in the future.
In our work, we propose to exploit the probabilistic nature of a quantum computer to sample multiple high-probability solutions of the balanced k-means problem at little additional cost. By formulating the k-means objective as a quadratic energy function, we embed the clustering task into the quantum-physical system of an AQC. By repeatedly measuring the quantum system, we sample high-probability solutions according to the Boltzmann distribution. Unlike previous approaches that only use the best solution and discard all other measurements, we utilize all samples to generate probabilistic solutions for the k-means problem, as shown in Figure 1. We recalibrate the samples of the clustering problem to address temperature mismatch (Pochart et al., 2021), estimating the posterior probability for each solution, which allows us to identify ambiguous points and provides alternative solutions. We demonstrate the algorithm on a D-Wave quantum computer and perform extensive experiments in simulation. However, our primary objective is to explore the potential and efficient utilization of quantum computing for machine learning. The primary contributions of our work are as follows:
  • We propose a quantum computing formulation of balanced k-means clustering that predicts well-calibrated confidence values and provides a set of alternative clustering solutions.
  • A reparametrization approach is used to compute posterior probabilities from samples that avoids exact tuning of the AQC sampling temperature.
  • Extensive experiments on synthetic and real data show the calibration of our approach using simulation as well as the D-Wave Advantage 2 AQC prototype.
2 Related Work
With the availability of quantum computers to the general research community (Bunyk et al., 2014; McGeoch & Farré, 2021; McGeoch et al., 2022; Debnath et al., 2016), the research interest in finding applications for such systems has considerably increased. In this context AQC (Kadowaki & Nishimori, 1998; Bunyk et al., 2014) provides a well-tangible starting point, even though many applications need a complete reformulation considering the architectural differences of a quantum computer. Current applications for AQC include solving optimization problems for improved traffic flow (Neukart et al., 2017), robotic routing problems (Ohzeki et al., 2019), the design of molecules (Mulligan et al., 2020) or portfolio optimization in the finance sector (Mugel et al., 2021). Recently, the computer vision community has developed a strong interest in finding applications for quantum computing, which are related to hard permutation problems (Birdal et al., 2021) like tracking (Zaech et al., 2022) and graph-matching (Benkner et al., 2021).
Clustering is a well-studied problem for quantum- as well as quantum-inspired algorithms (Aïmeur et al., 2007). Quantum clustering (Horn & Gottlieb, 2001) uses the Schrödinger equation to model the clustering problem, where cluster centers are defined as the minima of the corresponding potential function. Casaña-Eslava et al. (2020) extend this formulation with a probabilistic estimate of cluster memberships. These approaches use classical computation and gate-based quantum computers that are currently still on a small scale.
Closest to this work, Arthur & Date (2021) present a balanced k-means clustering algorithm suitable for an AQC. While our underlying clustering algorithm follows the same approach, Arthur & Date (2021) discard all but the best measurement. Our approach, however, utilizes all information by employing AQC as a sampler to generate probabilistic solutions of the clustering problem, rather than only using the best solution in an optimization framework.
3 Theory
3.1 Quantum Computing
Quantum computing is a fundamentally new approach, that utilizes the state of a quantum system to perform computations. In contrast to a classical computer, the state is probabilistic and described by its wave function, which enables the use fundamental properties of quantum systems like superposition and entanglement. Quantum algorithms that can efficiently run on these systems allow to solve previously infeasible problems, including the well-known prime factorization by Shor’s algorithm (Shor, 1999). In the following, the most important basics of quantum computing are explained.
Qubit.
The qubit forms the basic building block of a quantum computer. In contrast to the classical bit, which is always in either of its two basis states |$\ket{0},\ket{1}$, the qubit can exist in a state of superposition, which is a linear combination of any two basis states
|$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}.$ (1)
The two complex-valued factors $\alpha,\beta$ describing the superposition are called amplitudes.
Entanglement.
To perform meaningful operations, a quantum computer operates on a system of multiple qubits. If the qubits are independent of each other, the joint state of the system can be computed as the tensor-product of all involved qubit states. If the qubits of a system are entangled (Einstein et al., 1935; Schrödinger, 1935), it is is not possible to decompose the joint state into the tensor-product of each separate qubit state. Therefore, measuring the state of one qubit in a entangled system, affects the state of the other qubits. This is a fundamental property of quantum mechanics and plays a crucial role for the capabilities of quantum computing.
Measurement.
During computation on the quantum computer, the state of the system can be any valid superposition of basis states. Nevertheless a measurement of the system always results in a single basis state, which is referred to as wave-function collapse (Wilde, 2017). The probability of measuring a state is the respective squared amplitude. In the single qubit case as described in equation 1, this corresponds to
$p(\ket{0})=|\alpha|^{2}\quad\quad p(\ket{1})=|\beta|^{2}.$ (2)
In contrast to all other operations in quantum computing, this is not reversible.
3.2 Adiabatic Quantum Computing.
AQC, which is used in this work, is a quantum computing paradigm where the state of a quantum system is modified by performing an adiabatic transition. Current hardware implementations such as the D-wave systems (Bunyk et al., 2014) follow this approach and implement a Quantum Annealer (QA) to solve quadratic unconstrained binary optimization (QUBO). They are based on the Ising model (Ising, 1925; Kadowaki & Nishimori, 1998), which describes the configuration of a set of interacting particles that all carry an atomic spin $\sigma_{i}$.
The spin can either be +1 or -1 and the particles are coupled by interactions $J_{ij}$ as well as influenced individually by a transversal magnetic field $h_{i}$. The energy of this system is described by its Hamiltonian function
$H(\mathbf{\sigma})=-\sum_{i}\sum_{j}J_{ij}\sigma_{i}\sigma_{j}-\sum_{i}h_{i}\sigma_{i}.$ (3)
Thus, finding the lowest energy state or ground state of the Ising model corresponds to solving the QUBO defined by the Hamiltonian function. This relation is used in the QA, where the system of qubits implements an Ising model $H_{T}$ that represents the QUBO of interest.
The lowest energy state is found by following the adiabatic theorem (Born & Fock, 1928). Starting with an initial system in its ground state described by an Hamiltonian $H_{0}$, the adiabatic theorem states that during a sufficiently slow change of the Hamiltonian the system never leaves its ground state. The change of the Hamiltonian
$H(t)=H_{0}(1-\frac{t}{T_{a}})+H_{T}\frac{t}{T_{a}}$ (4)
is called an adiabatic transition. The transition is performed over the annealing-time $T_{a}$ and allows to solve an optimization problem with the Hamiltonian $H_{T}$.
While in an ideal noise-free case, the system stays in its ground state, any real system is embedded in a temperature bath that can induce a change to a higher energy state. The distribution of measured final states will then follow the Boltzmann distribution with temperature $T$
$p(\mathbf{\sigma})=\frac{\mbox{exp}[\nicefrac{{-H(\mathbf{\sigma})}}{{T}}]}{\sum_{\mathbf{\sigma}^{\prime}}\mbox{exp}[\nicefrac{{-H(\mathbf{\sigma}^{\prime})}}{{T}}]},$ (5)
where $p(\mathbf{\sigma})$ describes the probability of finding the system in state $\mathbf{\sigma}$ and $\mathbf{\sigma}^{\prime}$ are all possible states. In this work, we use this property, as it allows us to sample from the Boltzmann distribution corresponding to the energy-based model of clustering.
3.3 Energy-Based Models
Energy-based models are probabilistic models that use an energy function $E(\mathbf{x})$ to describe the probability of each system state. By assigning an energy value to each system state, the probabilities are Boltzmann distributed
$p(\mathbf{x})=\frac{1}{A}\mbox{exp}[\nicefrac{{E(\mathbf{x})}}{{T}}],$ (6)
with the temperature $T$ as a free parameter in this model and the normalization constant $A$ ensuring that the values are valid probabilities. Our model follows this approach and represents the clustering problem as an energy-based formulation.
One challenging task to use energy-based models remains in sampling the system according to the Boltzmann distribution as this is NP-hard for the quadratic Ising-model formulation (Pochart et al., 2021). While this is usually approximated on classical hardware with the corresponding computational overhead, AQC provides a direct way to sample from a physical system that follows the Boltzmann distribution (Pochart et al., 2021).
3.4 Clustering
Clustering is one core task required in many unsupervised machine learning algorithms. It has the objective to group a set of points $\mathbf{X}$ into disjoint clusters $\{c_{1},...,c_{K}\}$, where each cluster contains points that are similar in the feature space. A design parameter is the definition of the similarity metric, which leads to a wide range of clustering algorithms available. The popular and simple k-means algorithm (Forgy, 1965; Lloyd, 1982) optimizes the quadratic distance to a centroid. Such distance-based metrics assume clusters to be compact, while other approaches such as dbscan (Ester et al., 1996) define features as similar if they come from a contiguous region of high density.
In this work, balanced clustering is investigated, where a defined target cluster size $s_{k}$ is required. This approach forms the basis of many AQC-based algorithms (Benkner et al., 2020; Birdal et al., 2021; Zaech et al., 2022) and has a wide range of further applications such as secret key generation in cryptography (Hong et al., 2017), energy-efficient data aggregation in wireless sensor networks (Liao et al., 2013), or data cleaning (Fisher et al., 2015).
Uncertainty estimation in clustering and machine learning aims at quantifying the confidence that predictions made by a model correspond to the underlying ground-truth. The confidence scores can be used to eliminate uncertain samples from the solution, to select a set of possible hypotheses or to find the right parameters of the algorithm itself (Casaña-Eslava et al., 2019).
Uncertainty estimates can be generated separately for each data-point by evaluating the likelihood function $p(\mathbf{X}|{\hat{Z}})$, or by computing the posterior probability $p({\hat{Z}}|\mathbf{X})$ that describes the probability of the whole clustering solution. As the former can be evaluated if the data generating model is defined, it can be combined with most clustering approaches. Nevertheless, as the likelihood function is not a probability distribution of the cluster assignments, it does not provide a calibrated prediction. In contrast to this, the posterior probability $p({\mathbf{Z}}|\mathbf{X})$ directly represents the probability that the estimated assignment $\mathbf{Z}$ corresponds to the ground-truth. While this provides an interpretable result, it is often infeasible to compute as all possible assignments need to be evaluated in order to find the normalizing constant.
The set of possible clustering solutions together with their posterior probabilities provides valuable information that can be utilized for a range of low- as well as high-level reasoning tasks. On a low level, the probability of the best solutions can be used to identify the right number of clusters. If a number different than the number of the data-generating process is chosen, additional ambiguity is introduced. For a larger number during clustering than present in the data, the strong overlap between the corresponding distributions induces high ambiguity, while an insufficient number of clusters spreads points inside a single cluster far apart, which reduces the associated probability.
For high-level applications, calibrated samples of possible clustering solutions can provide additional information in multi-object tracking (Chiu et al., 2021). Following the AQC framework, tracking can be implemented as a clustering problem with additional constraints that represent the temporal relation between points (Zaech et al., 2022). Each cluster then corresponds to one object in the video, with each point representing it at a different timestep. The feature used to define clusters can either be visual similarity from re-identification(Hirzer et al., 2011; Paisitkriangkrai et al., 2015; Zheng et al., 2019) features or spatial similarity(Bewley et al., 2016). Different solutions thus represent different tracks through time, where ambiguities can be generated by occlusions or crossing paths. In a larger system such as an autonomous vehicle, knowing all likely candidate solutions can allow for predicting multimodal candidates for future trajectories (Jiang et al., 2023; Deo & Trivedi, 2018) and can help to evaluate the risk of taking any action. In a similar approach, feature matching can be formulated (Birdal et al., 2021) and ambiguous candidates can be discarded during 3d reconstruction or camera pose estimation.
3.5 Clustering as QUBO
To solve the clustering problem using AQC, a QUBO formulation is required. We use a variation of the one-hot encoding approach (Date et al., 2021; Arthur & Date, 2021) that uses a matrix ${Z}\in\{0,1\}^{K\times I}$ to encode the cluster assignment of $I$ samples to $K$ clusters. Each row corresponds to one of $K$ clusters and each column to one of $I$ samples. An entry ${Z}_{ki}=1$ indicates that the sample $x_{i}$ belongs to cluster $c_{k}$.
As each sample needs to be assigned to a single cluster, the sum of each column of $Z$ needs to satisfy the constraint $\sum_{k}{Z}_{ki}=1~{}\forall i$. The implementation of constrained clustering, where each cluster has a fixed size $s_{k}$ furthermore requires the row constraints $\sum_{i}{Z}_{ki}=s_{k}~{}\forall k$ on $Z$.
With the cost $q(i,j,k)$ for assigning the pair of samples $x_{i}$ and $x_{j}$ to the same cluster $c_{k}$, the optimization problem reads
$\hat{{Z}}=\operatorname*{arg\,min}_{Z}\sum_{k}\sum_{i}\sum_{j}Z_{ki}Z_{kj}q(i,j,k)$ (7)
[…]
By vectorizing $Z$ in row-major order as $\mathbf{z}=\text{vec}(Z)$, the optimization problem can be rewritten in matrix form as
$\hat{\mathbf{z}}=\operatorname*{arg\,min}_{\mathbf{z}}\mathbf{z}^{\intercal}\,\mathbf{Q}\,\mathbf{z}\quad\text{s.t.}\quad\mathbf{G}\mathbf{z}=\mathbf{d},$ (8)
where $\mathbf{Q}$ is a block diagonal matrix with blocks $\mathbf{Q}_{0},...,\mathbf{Q}_{K}$ and $\mathbf{G}\mathbf{z}=\mathbf{d}$ corresponds to the matrix formulation of the constraints. Each block $\mathbf{Q}_{k}$ of the cost matrix is a square matrix that contains the costs $q(i,j,k)$ at $\mathbf{Q}_{k,ij}$.
While the constraints are not directly covered by the QUBO problem, they can be modeled using Lagrangian multipliers. To avoid a mixed discrete and continuous optimization problem, a quadratic penalty reformulation $\lambda||\mathbf{G}\mathbf{x}-\mathbf{d}||_{2}^{2}$ is chosen, leading to the minimization problem
$\hat{\mathbf{z}}=\operatorname*{arg\,min}_{\mathbf{z}}\mathbf{z}^{\intercal}\mathbf{Q^{\prime}}\mathbf{z}+\mathbf{b^{\prime}}^{\intercal}\mathbf{z}$ (9)
with
$\mathbf{Q^{\prime}}=\mathbf{Q}+\lambda\mathbf{G}^{\intercal}\mathbf{G}$ (10)
$\mathbf{b^{\prime}}=-2\lambda\mathbf{G}^{\intercal}\mathbf{b}.$ (11)
In contrast to a linear penalty approach, where the multiplier $\lambda$ needs to be optimized, our selection only requires a sufficiently high $\lambda$, as all constraint violations result in an increased penalty term. With such selection, the penalty term evaluates to zero $\lambda||\mathbf{G}\mathbf{x}-\mathbf{d}||_{2}^{2}=0$ if all constraints are fulfilled, and thus, the minimizer $\hat{\mathbf{z}}$ of the modified optimization problem is a minimizer of the original optimization problem in Equation equation 8.
Besides modeling balanced clustering, further constraints can be introduced using Lagrangian multipliers. This allows to represent a wide range of tasks as QUBO clustering problems solvable with AQC.
4 Probabilistic Quantum Clustering
4.1 Motivation
Our clustering formulation employs a mixture of Gaussians to explain the observations, where the cluster assignments are latent variables. Each sample in a cluster $c_{k}$ is modeled as a sample drawn from a Gaussian distribution $\mathcal{N}(\mathbf{\mu}_{k},\mathbf{I})$ with mean $\mathbf{\mu}_{k}$ and identity covariance according to the k-means setting. Following a Bayesian approach, the best clustering solution can be found by maximizing the posterior probability over possible assignments $Z$ of points to clusters.
$p(Z|\mathbf{X})=\frac{p(\mathbf{X}|{Z})p({Z})}{\sum_{{Z}^{\prime}}p(\mathbf{X}|{Z^{\prime}})}=\frac{p(\mathbf{X}|{Z})p({Z})}{A}.$ (12)
While this approach provides a probabilistic estimate by jointly modeling information about the possible cluster configurations and the distribution of data points, it is often intractable to evaluate due to the partitioning function $A=\sum_{{Z}^{\prime}}p(\mathbf{X}|{Z^{\prime}})$. Computing $A$ requires to sum over all possible solutions $Z$, which grows exponentially with the number of samples in the clustering problem. To overcome this, we utilize an AQC that samples directly from a Boltzmann distribution, which we parameterize according to the probabilistic clustering problem.
4.2 Data Model
Determining the cost function of the optimization problem is a design choice of the algorithm. For our probabilistic approach, a well-defined data distribution, that forms the basis of the cost function is required. While many tasks approached in quantum computer vision, such as tracking (Zaech et al., 2022) or synchronization (Benkner et al., 2021; Birdal et al., 2021), costs are based on learned metrics or heuristics, they can also be trained to reflect the properties required in our approach.
We therefore, follow the mixture of Gaussian model, where each cluster generates samples from a normal distribution. With the independence of observations and clusters, given the distribution parameters, the likelihood of the joint observations for an assignment $Z$ is given as the product of the individual likelihods
$f(\mathbf{X}|{Z})=\prod_{k=1}^{K}f(\mathbf{X}|{Z_{k}})=\prod_{k=1}^{K}\prod_{i\in Z_{k}}f(\mathbf{x}_{i}|{Z_{k}})=$ (13)
$\prod_{k=1}^{K}\prod_{i\in Z_{k}}\frac{1}{\sqrt{2\pi^{d}}}\mbox{exp}\left[-\frac{1}{2}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})^{\intercal}\mathbf{I}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})\right]=$ (14)
$\frac{1}{\sqrt{2\pi^{d}}}\mbox{exp}\left[-\sum_{k=1}^{K}\sum_{i\in Z_{k}}\frac{1}{2}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})^{\intercal}\mathbf{I}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})\right]$ (15)
where the likelihood $f(\mathbf{x}_{i}|{Z_{k}})$ corresponds to a Gaussian distribution that follows $\mathcal{N}(\mathbf{\mu}_{k},\mathbf{I})$ and $d$ is the dimensionality of the space. This result can be used to formulate the energy-based model with the energy function
$E(\mathbf{X}|Z)=\sum_{k=1}^{K}\sum_{i\in Z_{k}}\frac{1}{2}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})^{\intercal}\mathbf{I}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}}).$ (16)
Using the energy function to rewrite the posterior distribution leads to the Boltzmann distribution from Equation 5
$p(Z|\mathbf{X})=\frac{\mbox{exp}\left[-(E(\mathbf{X}|Z)+E(Z))\right]}{\sum_{{Z}^{\prime}}\mbox{exp}\left[-E(\mathbf{X}|Z^{\prime})\right]}.$ (17)
In this formulation, the energy of the assignment $E(Z)$ corresponds to the prior $p({Z})$, which models feasible and infeasible solutions by an indicator function. In the case of balanced clustering, this corresponds to allowing assignments that have each point assigned to exactly one cluster and a cluster size according to $s_{k}$.
Searching the most likely clustering solution corresponds to finding lowest energy solution on on the AQC, as well as to the Maximum-a-Posteriori $\hat{Z}_{\text{map}}$ estimate of the assignment
$\hat{Z}_{\text{map}}=\operatorname*{arg\,max}_{Z}p(Z|\mathbf{X})p({Z})=\operatorname*{arg\,min}_{Z}E(\mathbf{X}|Z)+E(Z).$ (18)
As the AQC qubit system is an Ising model, it requires formulating $E(\mathbf{X}|Z)$ and $E(Z)$ quadratic in the optimization variables $Z$, enabling the joint discovery of assignments and cluster means. This is achieved by using the maximum likelihood estimator of the mean, resulting in a quadratic energy formulation that fits the Ising model in Equation 9
$E_{k}(\mathbf{X}|Z)=\frac{1}{s_{k}}\sum_{i}\sum_{j}Z_{ki}Z_{kj}~{}(\mathbf{x}_{i}-\mathbf{x}_{j})^{\intercal}(\mathbf{x}_{i}-\mathbf{x}_{j}).$ (19)
The second energy term $E(Z)$ in Equation 17, modeling the prior distribution over possible assignments cannot exactly be embedded in the Ising model. We therefore use the quadratic penalty method as shows in Equation equation 9 to approximate the energy of an indicator function. Importantly, this does not influence the energy of feasible solutions, as $\lambda||\mathbf{G}\mathbf{x}-\mathbf{d}||_{2}^{2}$ evaluates to zero for all solutions that fulfill the clustering constraints.
4.3 Boltzmann Reparametrization
While ideally the measurements on AQC are direct samples from the Boltzmann distribution (D-Wave Systems Inc., 2017; Dixit et al., 2021), it requires solving a range of challenges that prohibit a direct use in our scenario (Pochart et al., 2021): Mapping between the cost function and the physical system implemented on the AQC requires Hamiltonian scaling (Pochart et al., 2021). Estimating the scaling factor is nontrivial (Nelson et al., 2022) and prohibitive for using samples directly in many cases. Additionally, hardware limitations including imperfection of the processor and the spin-bath polarization effect (Pochart et al., 2021) prevent the AQC to sample the Boltzmann distribution exactly. Finally, as the energy term $E(Z)$ can only be implemented using the penalty method, the sampling density is influenced by solutions not fulfilling the constraints.
We compensate for these limitations by evaluating the energy of all measured feasible solutions $Z^{\prime}$ and recompute $p(Z|\mathbf{X})$ by evaluating the partitioning function over these, which only requires sampling solutions at a sufficiently high temperature.
4.4 Maximum Pointsets
Having the set of the most likely clustering solutions $Z$ available together with their posterior probability $P(Z|\mathbf{X})$ allows to find an assignment $Z^{*}$ of a subset of points that solves the clustering problem with increased probability $P(Z^{*}|\mathbf{X})$. Such a set can be found by using Algorithm 1, which implements a greedy approach that disregards points that disagree between different clustering solutions.
1:$Z^{*}\leftarrow Z_{0}$
2:$p\leftarrow P(Z_{0}|\mathbf{X})$
3:$i\leftarrow 1$
4:while $p\leq p_{\text{min}}$ do
5:     $Z_{i}^{\prime}\leftarrow align(Z_{i},Z^{*})$
6$\triangleright$ Find the cluster permutation that minimizes the number of points assigned to different clusters in $Z_{i}$ and $Z^{*}$.
7:     $Z^{*}\leftarrow Z^{*}\cap Z_{i}^{\prime}$
8$\triangleright$ Remove points assigned to different clusters.
9:     $p\leftarrow p+P(Z_{i}|\mathbf{X})$
10:     $i\leftarrow i+1$
11:end while
12:return $Z^{*}$
Algorithm 1 MaxsetSearch
In the case of a sufficiently well sampled Boltzmann distribution, the resulting maxset assignment $Z^{*}$ with the corresponding probabilistic estimate $P(Z^{*}|\mathbf{X})$ is still calibrated.
4.5 Inference Parameter Optimization
Using the quadratic penalty method to implement the constraints requires finding suitable Lagrangian multipliers $\lambda$. Even though a very high multiplier theoretically guarantees finding a feasible solution, it also deteriorates the conditioning of the optimization problem. Therefore, a suitable Lagrangian multiplier lifts the cost of any constraint violation above all relevant solutions of the clustering problem, while keeping them low enough to avoid scaling the total energy of the problem up considerably. To estimate the multipliers, we follow an iterative procedure as also proposed in (Zaech et al., 2022).
In an initial step, balanced k-means (Malinen & Fränti, 2014) is used to find a feasible clustering solution. This solution is used to estimate Lagrangian multipliers that avoid any first-order violation of the constraints. In subsequent iterative optimization steps, the problem is solved in simulation, to find multipliers that result in a well conditioned problem.
Such optimization procedure is crucial due to the low fidelity of the current generation of AQCs, which requires careful engineering of the problem energy. Therefore, we expect this procedure to become of reduced importance with the progress of quantum computing.
5 Experiments and Results
We perform experiments on synthetic data as well as real data to verify the efficacy of our method in finding the set of high-probability solutions and in estimating calibrated confidence scores. The experimental scenarios are solved with QA, Simulated Annealing (SIM), and exact exhaustive search using the presented energy formulation and with k-means as a baseline method, which all optimize for the same cost objective. This further allows us to understand the limitations and required work when deploying the approach to real quantum computers.
5.1 Solver Methods
[…]
Due to the strong compute-time limitations on an AQC, all Lagrangian optimization steps are performed with SIM, before measuring the final results on the AQC.
Simulated annealing (SIM) provided by D-Wave is used for larger scale comparisons. Similar to QA, we perform 5000 runs for each clustering problem. We reduce the number of sweeps performed in each run to 30, which allows us to sample the Boltzmann distribution at a sufficiently high temperature in most scenarios, making it comparable to QA.
Exact exhaustive search is used as a reference method to validate the energy-based formulation on small problems. By iterating all feasible solutions, the lowest energy solution is guaranteed to be found and the partitioning function is computed exactly.
K-Means clustering with a balanced cluster constraint (Malinen & Fränti, 2014) forms the baseline for our approach. We run the algorithm until convergence for a maximum of 1000 iterations. While this solution does not provide a probabilistic estimate, it is useful to assess the relative clustering performance.
5.2 Dataset and Metrics
Data for the quantitative evaluation of our method is synthetically generated. For each clustering problem a total of $I$ points are sampled from a separate normal distribution for each of $K$ clusters. The cluster centers are randomly drawn, such that the distance between each pair of clusters lies within a predefined range $[d_{\text{min}},d_{\text{max}}]$. For each experiment a total of $L$ clustering tasks is generated. This allows us to evaluate the calibration metrics over a large value range. For all experiments that directly compare methods, identical clustering tasks are used.
Further qualitative examples are provided for the IRIS dataset (Fisher, 1936), which contains 50 samples of 4 features in 3 classes. We randomly subsample the points and dimensions to generate the parameters required for our experiments.
Clustering metrics are computed using the available ground-truth clusters. We evaluate 4 standard metrics: The accuracy, which measures the ratio of clustering solutions that are identical to the ground-truth. Completeness (Rosenberg & Hirschberg, 2007) measures the ratio of points from a single cluster being grouped together. The adjusted Rand score (Hubert & Arabie, 1985) compares all pairs of points in the ground truth and prediction and the Fowlkes-Mallows index (Fowlkes & Mallows, 1983) combines precision and recall into a single score.
Solver
Method
Accuracy
Completeness
Adjusted
Rand Index
Fowlkes-
Mallows Score
Accuracy
Completeness
Adjusted
Rand Index
Fowlkes-
Mallows Score
Accuracy
Completeness
Adjusted
Rand Index
Fowlkes-
Mallows Score
15 Points, 3 Clusters, 2 Dim 30 Points, 3 Clusters, 2 Dim 45 Points, 3 Clusters, 2 Dim
SIM 56.4$\pm$1.6 79.5$\pm$0.8 74.7$\pm$1.0 81.9$\pm$0.7 38.6$\pm$1.5 74.2$\pm$0.8 73.3$\pm$0.9 81.6$\pm$0.6 50.8$\pm$1.6 86.2$\pm$0.5 87.3$\pm$0.5 91.4$\pm$0.3
K-means 51.3$\pm$1.6 75.0$\pm$0.9 68.8$\pm$1.1 77.7$\pm$0.8 37.0$\pm$1.5 71.9$\pm$0.8 70.2$\pm$0.9 79.4$\pm$0.6 52.8$\pm$1.1 85.9$\pm$0.4 86.6$\pm$0.4 90.9$\pm$0.3
15 Points, 3 Clusters, 2 Dim 10 Points, 2 Clusters, 2 Dim 20 Points, 4 Clusters, 4 Dim
QA 56.1$\pm$1.6 79.4$\pm$0.8 74.6$\pm$1.0 81.9$\pm$0.7 74.3$\pm$1.4 80.2$\pm$1.1 79.6$\pm$1.1 88.7$\pm$0.6 12.4$\pm$1.0 51.2$\pm$0.8 34.0$\pm$1.0 47.9$\pm$0.8
SIM 56.4$\pm$1.6 79.5$\pm$0.8 74.7$\pm$1.0 81.9$\pm$0.7 74.1$\pm$1.4 80.0$\pm$1.1 79.5$\pm$1.1 88.6$\pm$0.6 32.4$\pm$1.5 70.4$\pm$0.8 59.2$\pm$1.1 67.8$\pm$0.8
K-means 51.3$\pm$1.6 75.0$\pm$0.9 68.8$\pm$1.1 77.7$\pm$0.8 70.1$\pm$1.4 76.3$\pm$1.2 75.4$\pm$1.2 86.3$\pm$0.7 20.9$\pm$1.3 61.9$\pm$0.8 47.4$\pm$1.0 58.5$\pm$0.8
Exhaustive 56.4$\pm$1.6 79.5$\pm$0.8 74.7$\pm$1.0 81.9$\pm$0.7 74.3$\pm$1.4 80.2$\pm$1.1 79.6$\pm$1.1 88.7$\pm$0.6 - - - -
Table 1: Synthetic data results for our approach (QA, SIM, exhaustive) and k-means. All numbers in % with standard error of the mean.
5.3 Calibration Performance
We evaluate the calibration of our method on a synthetic clustering scenario with 3 clusters and 15 points using QA, SIM and exhaustive search on 1000 tasks. First all clustering solutions $Z$ are accumulated in bins according to their estimated posterior probability $P(Z|\mathbf{X})$. This process also includes all sampled non-optimal but feasible solutions. The resulting histogram of solutions is shown in Figure 1(a). After accumulation, the ratio of correct solutions in each bin is evaluated in Figures 1(b),1(c), and 1(d), which for a calibrated method should be close to the mean predicted probability represented by the diagonal. We find our approach to generate well-calibrated probabilities in both simulation and when using QA.
(a) Histogram of predicted probabilities.
(b) 2 clusters, 10 points.
(c) 3 clusters, 15 points.
(d) 4 clusters, 20 points.
Figure 2: Evaluation of the calibration for QA, SIM and exhaustive search in a clustering scenario with 15 points and 3 clusters. Evaluation of the calibration for quantum annealing, simulated annealing and exhaustive search in clustering scenarios with 2, 3, and 4 clusters and 5 points in each cluster. All results are generated with 1,000 problems in each scenario and 5,000 measurements for each clustering problem.
5.4 Clustering Performance
We evaluate the performance of our clustering formulation using synthetic data in Table 1. The upper rows show results for 3 clusters with an increasing number of points and 1000 tasks/setting for SIM and k-means. Due to the problem size, results cannot be found using QA and exhaustive search. Comparing results shows that most clustering metrics are better when using our formulation with SIM, however, the difference decreases with increasing problem size. This can be attributed to the globally optimal solution our approach is optimizing for. While in an ideal the global solution is at least as good as the k-means solution, an increasing problem size also increases the complexity of the corresponding QUBO. This makes it harder to solve exactly and the noisy approach using annealing is not capable of solving the problem correctly.
For the largest clustering problem with 3 clusters and 45 points, k-means provides the best accuracy and thus, a larger number of correct solutions compared to SIM. However, the metrics indicating clustering quality are higher for SIM. This indicates that our formulation is able to find a better solution in cases where the problem is not solved correctly by SIM and k-means.
Figure 3: Sparsification plots of clustering metrics.
The lower rows in Table 1 show settings with 2,3, and 4 clusters, each containing 5 points, again with 1000 tasks/setting. For the two smaller scenarios QA on the D-wave Advantage 2 provides results close or identical to SIM and exhaustive search. For the largest scenario with 20 points in 4 clusters, QA loses performance.
The quality of predicted posterior probabilities can be further evaluated by linking them to the clustering metrics and sparsifying the set of tasks. Starting with metrics over the whole set of clustering tasks, we remove tasks according to increasing predicted probability and evaluate the metric over the remaining set. In an ideal predictor, this removes the lowest performing tasks first. In Figures LABEL:fig:sparsification_method_comp and LABEL:fig:sparsification_metric_comp this is done for the adjusted Rand index with different solvers and for QA with different metrics respectively. The solid lines show the sparsification-plots using the predicted probabilities and the dashed lines show the same plots for an oracle method that generates the best possible ordering based on the metrics themselves.
In Figure LABEL:fig:sparsification_method_comp it becomes apparent that QA, SIM and exhaustive search perform close to each other over most of the value range. Nevertheless, for a high sparsification with more than 80% of the tasks removed, QA shows a drop in performance compared to the other methods. This is caused by tasks where only a single, but incorrect solution is found, which gets assigned a posterior probability of $P(Z|\mathbf{X})=1.0$. In such cases the Boltzmann distribution has not been sampled sufficiently well, either because of too few measurements or because of a low effective sampling temperature. As SIM does not show this behavior the source can likely be traced back to current limitation of the quantum computer.
5.5 Maximum Pointsets
Qualitative examples for the maximum pointsets generated with Algorithm 1 are depicted in Figure 4. The shape of each point represents the ground truth class and the color the assigned cluster. Starting with the most likely solution having a predicted probability of $p(Y|X)=0.61$ on the left, each plot shows one additional step of the algorithm, which successively removes points from the solution, indicated by plotting them in Grey. The illustration demonstrates that the MaxsetSearch algorithm is able to generate well-separated clusters from the probabilistic predictions.
(a) $p=0.61$
(b) $p=0.77$
(c) $p=0.86$
(d) $p=0.90$
Figure 4: Visualization of max pointsets for synthetic data with the probability for each of the determined pointsets.
5.6 IRIS Dataset
While our method assumes an identity covariance in the data, it can be applied to other distributions, which we evaluate on the widely used IRIS dataset. Clustering metrics for experiments using 3 clusters with 5 points each are provided in Table 6 and show that our formulation using QA and SIM is competitive with k-means. Figures LABEL:fig:qual_iris_qa and LABEL:fig:qual_iris_sim show differently sized qualitative examples from the dataset solved using our formulation with QA and SIM respectively.
[…]
Figure 5: Qualitative results on the IRIS dataset.
Figure 6: Clustering metrics generated on IRIS subsets using 3 clusters with 15 points in total.
Though the value of the predicted probability cannot be interpreted easily with the mismatch between the assumed and the actual data-generating process, it can be used to find maximum pointsets. The results from Algorithm 1 using results from SIM are provided in Figure 7 and show that even for a distribution mismatch the maximum pointsets can provide meaningful results for removing ambiguous samples.
Figure 7: Max pointsets generated using simulated annealing on the IRIS dataset with 60 points.
6 Conclusion
In this work, we proposed a probabilistic clustering approach based on sampling k-means solutions using AQC. By using all valid measurements, calibrated confidence scores are computed at little cost and solutions are competitive to an iterative balanced k-means approach. We evaluated our method on synthetic as well as real data using simulation, exhaustive search as well as the D-Wave Advantage 2 prototype AQC to explore the potential of quantum computing in machine learning and computer vision.
While the approach is still limited in the problem size, quantum computing enables a fundamentally different approach to clustering that can provide additional information that is costly to compute otherwise. Nevertheless, even with the current progress and potential to scale to real-world problems, more work is required to adapt existing problem formulations, such that the full capability of quantum computing can be effectively used.
Acknowledgments
This work was funded by Toyota Motor Europe via the research project TRACE Zürich.

### v2

@@
-Probabilistic Sampling of Balanced K-Means using Adiabatic Quantum Computing
+heightadjust=object
 Abstract
-Adiabatic quantum computing (AQC) is a promising quantum computing approach for discrete and often NP-hard optimization problems.
+Adiabatic quantum computing (AQC) is a promising approach for discrete and often NP-hard optimization problems.
 Current AQCs allow to implement problems of research interest, which has sparked the development of quantum representations for many computer vision tasks.
@@
 Instead of discarding non-optimal solutions, we propose to use them to compute calibrated posterior probabilities with little additional compute cost.
-This allows us to identify ambiguous solutions and data points, which we demonstrate on a D-Wave AQC on synthetic and real data.
+This allows us to identify ambiguous solutions and data points, which we demonstrate on a D-Wave AQC on synthetic tasks and real visual data.
+1 Introduction
 Figure 1: The proposed approach uses an adiabatic quantum computer to sample solutions of a balanced k-means problem.
@@
 By reparametrizing the distribution, the calibrated posterior probability of each solution can be estimated.
-1 Introduction
-Clustering is a fundamental problem in machine learning and computer vision, extensively employed for the analysis and organization of large volumes of unlabeled visual data.
-It involves grouping similar objects based on their features, facilitating efficient data processing and unsupervised learning algorithms (Xu & Wunsch, 2005).
-The information extracted from clusters plays a vital role in various computer vision applications, including image classification (Lazebnik et al., 2006; Caron et al., 2018), segmentation (Coleman & Andrews, 1979; Dhanachandra et al., 2015), tracking (Keuper et al., 2020), and network training (Coates & Ng, 2012; Yang et al., 2016).
-However, the often NP-hard computational complexity of solving clustering problems often hinders their application to large-scale problems that require fast processing times (Xu & Wunsch, 2005).
-In addition to this, the ambiguity of the task itself contributes to the absence of a unique definition for clustering (Xu & Wunsch, 2005), which arises from the range of possible optimization objectives.
-An example of this are different algorithms that can optimize for the most compact solution, an underlying data-generating process or just use local distances.
-One way to address this challenge is the use of confidence-based estimators and the sampling of multiple likely solutions.
-This approach helps in identifying well-assigned samples and uncovering ambiguities within the data (Xu & Wunsch, 2005).
-Quantum algorithms that promise a considerable speedup over their classical counterparts and are inherently probabilistic have the potential to enable a new family of machine learning algorithms.
-By now, quantum machine learning algorithms approach tasks such as optimization of quadratic problems (Kadowaki & Nishimori, 1998), training of restricted Boltzmann machines (Dixit et al., 2021), and learning with quantum neural networks (Abbas et al., 2021).
+Clustering and uncertainty quantification (UQ) are fundamental problems in machine learning and computer vision.
+Clustering, the task of grouping objects based on the similarity of their features, plays a pivotal role in analysis and organization of vast quantities of unlabeled visual data with applications including image classification [51, 16, 65], segmentation [23, 32], tracking [48], and network training [75, 22, 76].
+UQ, on the other hand, aims to assess the trustworthiness of predictions and accounts for the impact of data variability.
+As a crucial element for certifying confidence in results, UQ is indispensable in clustering, especially given the often ambiguous nature of data partitioning [41].
+Clustering typically involves solving for an optimal assignment of data points to (latent) centroids, while a full Bayesian UQ approach strives to accurately depict the posterior distribution.
+Unfortunately, both of these problems are known to be notoriously difficult, classified within the NP complexity class [^1] Computing marginal probabilities on a Bayesian network is, basically, a counting problem that can become arbitrarily hard to solve.
+[24, 26, 55].
+These challenges within the Turingian modes of computation have motivated: (i) continuous relaxations in clustering [78] and (ii) the advent of approximate Bayesian inference methods for UQ, including the EM algorithm [30], variational inference [13], and sequential MCMC [29].
+In this paper, departing from the above-mentioned approximations or continuous relaxations, we take a completely different approach and harness the upcoming and novel computational paradigm of quantum computers to tackle the challenging task of clustering with calibrated uncertainty quantification.
+Specifically, we focus on the widely used balanced K-means [52, 54, 56], which is an iterative algorithm that first assigns each data point to a cluster followed by an update of cluster centroids.
+Despite their early state, quantum machine learning algorithms have approached tasks such as optimization of quadratic problems [47], training of restricted Boltzmann machines [33], and learning with quantum neural networks [1].
 While the current quantum computers can only solve small-scale problems, they provide the basis to develop and test algorithms that can considerably increase the size of feasible problems in the future.
-In our work, we propose to exploit the probabilistic nature of a quantum computer to sample multiple high-probability solutions of the balanced k-means problem at little additional cost.
-By formulating the k-means objective as a quadratic energy function, we embed the clustering task into the quantum-physical system of an AQC.
+In our work, we interpret K-means clustering in an energy-based framework and aim to model the posterior distribution by sampling multiple high-probability binary assignments from it on actual quantum hardware at little additional cost.
+We specifically use the quantum annealer (QA) of D-Wave [58], which follows the concept of adiabatic quantum computing (AQC) [10] and avoids simulating costly Markov chains as in classical computers.
+To achieve this, we first formulate the k-means objective as a quadratic energy function over binary variables and embed the clustering task into the quantum-physical system of D-Wave.
 By repeatedly measuring the quantum system, we sample high-probability solutions according to the Boltzmann distribution.
 Unlike previous approaches that only use the best solution and discard all other measurements, we utilize all samples to generate probabilistic solutions for the k-means problem, as shown in Figure 1.
-We recalibrate the samples of the clustering problem to address temperature mismatch (Pochart et al., 2021), estimating the posterior probability for each solution, which allows us to identify ambiguous points and provides alternative solutions.
-We demonstrate the algorithm on a D-Wave quantum computer and perform extensive experiments in simulation.
-However, our primary objective is to explore the potential and efficient utilization of quantum computing for machine learning.
-The primary contributions of our work are as follows:
-  • We propose a quantum computing formulation of balanced k-means clustering that predicts well-calibrated confidence values and provides a set of alternative clustering solutions.
-  • A reparametrization approach is used to compute posterior probabilities from samples that avoids exact tuning of the AQC sampling temperature.
-  • Extensive experiments on synthetic and real data show the calibration of our approach using simulation as well as the D-Wave Advantage 2 AQC prototype.
+Due to temperature mismatch [68], AQCs sample from a modified posterior instead of the true posterior, which needs to be re-calibrated for the task at hand.
+We do so by estimating the posterior probability for each solution.
+This allows us to identify ambiguous points and provide alternative solutions.
+By leveraging QA, our approach respects the binary nature of the assignment problem.
+We demonstrate our algorithm on a D-Wave quantum computer and also perform extensive experiments in simulation.
+On real data, we show that our approach can be utilized for distributions that do not strictly follow the initial assumptions and that it is suitable for identifying ambiguous images.
+Our primary contributions are:
+  • A quantum computing formulation of balanced k-means clustering that predicts calibrated confidence values and provides a set of alternative clustering solutions.
+  • A reparametrization approach used to calibrate posterior probabilities from samples that avoids exact tuning of the AQC sampling temperature.
+  • Extensive experiments on synthetic and real data showing the calibration of our approach both on simulations and on the D-Wave Advantage 2 QA prototype.
+Developing better heuristics for the NP-hard challenges of clustering and UQ presents significant difficulties for classical approaches.
+However, we are optimistic that the ongoing progress in quantum computing technology will progressively enhance the effectiveness of our algorithms.
 2 Related Work
-With the availability of quantum computers to the general research community (Bunyk et al., 2014; McGeoch & Farré, 2021; McGeoch et al., 2022; Debnath et al., 2016), the research interest in finding applications for such systems has considerably increased.
-In this context AQC (Kadowaki & Nishimori, 1998; Bunyk et al., 2014) provides a well-tangible starting point, even though many applications need a complete reformulation considering the architectural differences of a quantum computer.
-Current applications for AQC include solving optimization problems for improved traffic flow (Neukart et al., 2017), robotic routing problems (Ohzeki et al., 2019), the design of molecules (Mulligan et al., 2020) or portfolio optimization in the finance sector (Mugel et al., 2021).
-Recently, the computer vision community has developed a strong interest in finding applications for quantum computing, which are related to hard permutation problems (Birdal et al., 2021) like tracking (Zaech et al., 2022) and graph-matching (Benkner et al., 2021).
-Clustering is a well-studied problem for quantum- as well as quantum-inspired algorithms (Aïmeur et al., 2007).
-Quantum clustering (Horn & Gottlieb, 2001) uses the Schrödinger equation to model the clustering problem, where cluster centers are defined as the minima of the corresponding potential function.
+Quantum computation.
+With the availability of quantum computers to the general research community [15, 57, 58, 28], the research interest in finding applications for such systems has considerably increased.
+In this context AQC [47, 15] provides a well-tangible starting point, even though many applications need a complete reformulation considering the architectural differences of a quantum computer.
+Current applications for AQC are optimization tasks in different fields [64, 66, 62, 61], including quantum computer vision [12], where a strong interest in finding quantum computing formulations has developed.
+These are often related to hard permutation problems [7, 12, 6] like tracking [77], graph, shape and point matching [6, 9, 60] or training binary (graph) neural networks [49].
+In this context, Birdal et al.
+[12] evaluate the k-best solutions, however, without the energy-based formulation, only little improvement is achieved.
+Another tasks of interest for the community is model fitting to find camera parameters [34] or separating motion components [3], which can be formulated as a quantum-computing problem [74, 20, 3, 34, 19, 37] instead of consensus maximization.
+Clustering.
+Clustering is a well-studied problem for quantum- as well as quantum-inspired algorithms [2].
+It groups a dataset ${X}$ into disjoint clusters ${c_{1},...,c_{K}}$, with each cluster containing points similar in the feature space.
+For example, k-means algorithm [39, 53] utilizes quadratic distances, favoring compactness, whereas dbscan [36] identifies local high-density regions.
+Quantum clustering methods use either the paradigm of AQC [5, 4, 65, 50] or circuit-based quantum computing such as Quantum clustering [43] that models clustering through the Schrödinger equation, where cluster centers are defined as the minima of its potential function.
 Casaña-Eslava et al.
-(2020) extend this formulation with a probabilistic estimate of cluster memberships.
-These approaches use classical computation and gate-based quantum computers that are currently still on a small scale.
-Closest to this work, Arthur & Date (2021) present a balanced k-means clustering algorithm suitable for an AQC.
-While our underlying clustering algorithm follows the same approach, Arthur & Date (2021) discard all but the best measurement.
+[18] extend this formulation with a probabilistic estimate of cluster memberships.
+These approaches use classical computation and circuit-based quantum computers that are fundamentally different from AQC and currently still on a smaller scale.
+Closest to this work, Arthur and Date [4] present a balanced k-means clustering algorithm with predefined target size $s_{k}$ suitable for an AQC and Nguyen et al.
+[65] use a similar formulation to cluster visual features.
+While our underlying algorithm follows a similar approach as [4, 65], they discard all but the best measurement.
 Our approach, however, utilizes all information by employing AQC as a sampler to generate probabilistic solutions of the clustering problem, rather than only using the best solution in an optimization framework.
-3 Theory
-3.1 Quantum Computing
+We particularly consider balanced clustering, as it forms the basis of several AQC algorithms [7, 12, 77] in computer vision.
+Uncertainty quantification in clustering.
+Knowing the set of high-quality solutions and their confidence offers valuable insights for both low- and high-level tasks.
+E.g.
+at a low level, the probability of top solutions can help to determine the correct cluster count.
+Choosing a different number than the actual data process introduces ambiguity [17]: too many clusters cause overlap, while too few clusters spread points, reducing associated probability.
+For high-level tasks, calibrated clustering solutions have the potential to enhance matching problems like multi-object tracking [21].
+Following the AQC framework, tracking becomes a clustering problem with additional constraints that represent the temporal relation between points [77], where each cluster corresponds to one track.
+The feature defining clusters can be visual similarities from re-identification [42, 67, 79] or spatial similarity [8].
+The set of solutions models different trajectories due to occlusions or intersecting paths.
+In complex systems like autonomous vehicles, understanding multiple probable solutions helps to predict multimodal future trajectories [46, 31], to assess risks when taking actions.
+In a similar approach, AQC is used for feature matching [12] and uncertainty estimation potentially allows to discard ambiguous candidates during 3d reconstruction.
+3 Adiabatic Quantum Computing
 Quantum computing is a fundamentally new approach, that utilizes the state of a quantum system to perform computations.
 In contrast to a classical computer, the state is probabilistic and described by its wave function, which enables the use fundamental properties of quantum systems like superposition and entanglement.
-Quantum algorithms that can efficiently run on these systems allow to solve previously infeasible problems, including the well-known prime factorization by Shor’s algorithm (Shor, 1999).
+Quantum algorithms that can efficiently run on these systems allow to solve previously infeasible problems, including the well-known prime factorization by Shor’s algorithm [71].
 In the following, the most important basics of quantum computing are explained.
 Qubit.
-The qubit forms the basic building block of a quantum computer.
-In contrast to the classical bit, which is always in either of its two basis states |$\ket{0},\ket{1}$, the qubit can exist in a state of superposition, which is a linear combination of any two basis states
-|$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}.$ (1)
+Qubits are the building block of a quantum computer.
+In contrast to the classical bit, which is in either of its two basis states |$\ket{0},\ket{1}$, the qubit can exist in a state of superposition, which is a linear combination of any two basis states |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$.
 The two complex-valued factors $\alpha,\beta$ describing the superposition are called amplitudes.
@@
 If the qubits are independent of each other, the joint state of the system can be computed as the tensor-product of all involved qubit states.
-If the qubits of a system are entangled (Einstein et al., 1935; Schrödinger, 1935), it is is not possible to decompose the joint state into the tensor-product of each separate qubit state.
+If the qubits of a system are entangled [35, 70], it is is not possible to decompose the joint state into the tensor-product of each separate qubit state.
 Therefore, measuring the state of one qubit in a entangled system, affects the state of the other qubits.
@@
 During computation on the quantum computer, the state of the system can be any valid superposition of basis states.
-Nevertheless a measurement of the system always results in a single basis state, which is referred to as wave-function collapse (Wilde, 2017).
+Nevertheless, a measurement of the system always results in a single state of the measurement basis, which is referred to as wave-function collapse [73].
 The probability of measuring a state is the respective squared amplitude.
-In the single qubit case as described in equation 1, this corresponds to
+In the single qubit case, this corresponds to
 $p(\ket{0})=|\alpha|^{2}\quad\quad p(\ket{1})=|\beta|^{2}.$ (1)
 In contrast to all other operations in quantum computing, this is not reversible.
-3.2 Adiabatic Quantum Computing.
+Adiabatic Quantum Computing.
 AQC, which is used in this work, is a quantum computing paradigm where the state of a quantum system is modified by performing an adiabatic transition.
-Current hardware implementations such as the D-wave systems (Bunyk et al., 2014) follow this approach and implement a Quantum Annealer (QA) to solve quadratic unconstrained binary optimization (QUBO).
-They are based on the Ising model (Ising, 1925; Kadowaki & Nishimori, 1998), which describes the configuration of a set of interacting particles that all carry an atomic spin $\sigma_{i}$.
+Current hardware implementations such as the D-wave systems [15] follow this approach by implementing QA to solve quadratic unconstrained binary optimization (QUBO).
+They are based on the Ising model [45, 47], which describes the configuration of a set of interacting particles that all carry an atomic spin $s_{i}$.
 The spin can either be +1 or -1 and the particles are coupled by interactions $J_{ij}$ as well as influenced individually by a transversal magnetic field $h_{i}$.
 The energy of this system is described by its Hamiltonian function
-$H(\mathbf{\sigma})=-\sum_{i}\sum_{j}J_{ij}\sigma_{i}\sigma_{j}-\sum_{i}h_{i}\sigma_{i}.$ (3)
+$H({s})=-\sum_{i}\sum_{j}J_{ij}s_{i}s_{j}-\sum_{i}h_{i}s_{i}.$ (2)
 Thus, finding the lowest energy state or ground state of the Ising model corresponds to solving the QUBO defined by the Hamiltonian function.
 This relation is used in the QA, where the system of qubits implements an Ising model $H_{T}$ that represents the QUBO of interest.
-The lowest energy state is found by following the adiabatic theorem (Born & Fock, 1928).
-Starting with an initial system in its ground state described by an Hamiltonian $H_{0}$, the adiabatic theorem states that during a sufficiently slow change of the Hamiltonian the system never leaves its ground state.
-The change of the Hamiltonian
+The lowest energy state is found by following the adiabatic theorem [14].
+Starting with an initial system in its ground state described by a Hamiltonian $H_{0}$, the adiabatic theorem states that during a sufficiently slow change of the Hamiltonian the system never leaves its ground state.
+The change of the Hamiltonian is called an adiabatic transition and is often performed with a linear schedule
 $H(t)=H_{0}(1-\frac{t}{T_{a}})+H_{T}\frac{t}{T_{a}}$ (3)
-is called an adiabatic transition. The transition is performed over the annealing-time $T_{a}$ and allows to solve an optimization problem with the Hamiltonian $H_{T}$.
+over the annealing-time $T_{a}$ and allows to solve an optimization problem with the Hamiltonian $H_{T}$.
 While in an ideal noise-free case, the system stays in its ground state, any real system is embedded in a temperature bath that can induce a change to a higher energy state.
 The distribution of measured final states will then follow the Boltzmann distribution with temperature $T$
-$p(\mathbf{\sigma})=\frac{\mbox{exp}[\nicefrac{{-H(\mathbf{\sigma})}}{{T}}]}{\sum_{\mathbf{\sigma}^{\prime}}\mbox{exp}[\nicefrac{{-H(\mathbf{\sigma}^{\prime})}}{{T}}]},$ (5)
-where $p(\mathbf{\sigma})$ describes the probability of finding the system in state $\mathbf{\sigma}$ and $\mathbf{\sigma}^{\prime}$ are all possible states.
-In this work, we use this property, as it allows us to sample from the Boltzmann distribution corresponding to the energy-based model of clustering.
-3.3 Energy-Based Models
-Energy-based models are probabilistic models that use an energy function $E(\mathbf{x})$ to describe the probability of each system state.
-By assigning an energy value to each system state, the probabilities are Boltzmann distributed
-$p(\mathbf{x})=\frac{1}{A}\mbox{exp}[\nicefrac{{E(\mathbf{x})}}{{T}}],$ (6)
-with the temperature $T$ as a free parameter in this model and the normalization constant $A$ ensuring that the values are valid probabilities.
-Our model follows this approach and represents the clustering problem as an energy-based formulation.
-One challenging task to use energy-based models remains in sampling the system according to the Boltzmann distribution as this is NP-hard for the quadratic Ising-model formulation (Pochart et al., 2021).
-While this is usually approximated on classical hardware with the corresponding computational overhead, AQC provides a direct way to sample from a physical system that follows the Boltzmann distribution (Pochart et al., 2021).
-3.4 Clustering
-Clustering is one core task required in many unsupervised machine learning algorithms.
-It has the objective to group a set of points $\mathbf{X}$ into disjoint clusters $\{c_{1},...,c_{K}\}$, where each cluster contains points that are similar in the feature space.
-A design parameter is the definition of the similarity metric, which leads to a wide range of clustering algorithms available.
-The popular and simple k-means algorithm (Forgy, 1965; Lloyd, 1982) optimizes the quadratic distance to a centroid.
-Such distance-based metrics assume clusters to be compact, while other approaches such as dbscan (Ester et al., 1996) define features as similar if they come from a contiguous region of high density.
-In this work, balanced clustering is investigated, where a defined target cluster size $s_{k}$ is required.
-This approach forms the basis of many AQC-based algorithms (Benkner et al., 2020; Birdal et al., 2021; Zaech et al., 2022) and has a wide range of further applications such as secret key generation in cryptography (Hong et al., 2017), energy-efficient data aggregation in wireless sensor networks (Liao et al., 2013), or data cleaning (Fisher et al., 2015).
-Uncertainty estimation in clustering and machine learning aims at quantifying the confidence that predictions made by a model correspond to the underlying ground-truth.
-The confidence scores can be used to eliminate uncertain samples from the solution, to select a set of possible hypotheses or to find the right parameters of the algorithm itself (Casaña-Eslava et al., 2019).
-Uncertainty estimates can be generated separately for each data-point by evaluating the likelihood function $p(\mathbf{X}|{\hat{Z}})$, or by computing the posterior probability $p({\hat{Z}}|\mathbf{X})$ that describes the probability of the whole clustering solution.
-As the former can be evaluated if the data generating model is defined, it can be combined with most clustering approaches.
-Nevertheless, as the likelihood function is not a probability distribution of the cluster assignments, it does not provide a calibrated prediction.
-In contrast to this, the posterior probability $p({\mathbf{Z}}|\mathbf{X})$ directly represents the probability that the estimated assignment $\mathbf{Z}$ corresponds to the ground-truth.
-While this provides an interpretable result, it is often infeasible to compute as all possible assignments need to be evaluated in order to find the normalizing constant.
-The set of possible clustering solutions together with their posterior probabilities provides valuable information that can be utilized for a range of low- as well as high-level reasoning tasks.
-On a low level, the probability of the best solutions can be used to identify the right number of clusters.
-If a number different than the number of the data-generating process is chosen, additional ambiguity is introduced.
-For a larger number during clustering than present in the data, the strong overlap between the corresponding distributions induces high ambiguity, while an insufficient number of clusters spreads points inside a single cluster far apart, which reduces the associated probability.
-For high-level applications, calibrated samples of possible clustering solutions can provide additional information in multi-object tracking (Chiu et al., 2021).
-Following the AQC framework, tracking can be implemented as a clustering problem with additional constraints that represent the temporal relation between points (Zaech et al., 2022).
-Each cluster then corresponds to one object in the video, with each point representing it at a different timestep.
-The feature used to define clusters can either be visual similarity from re-identification(Hirzer et al., 2011; Paisitkriangkrai et al., 2015; Zheng et al., 2019) features or spatial similarity(Bewley et al., 2016).
-Different solutions thus represent different tracks through time, where ambiguities can be generated by occlusions or crossing paths.
-In a larger system such as an autonomous vehicle, knowing all likely candidate solutions can allow for predicting multimodal candidates for future trajectories (Jiang et al., 2023; Deo & Trivedi, 2018) and can help to evaluate the risk of taking any action.
-In a similar approach, feature matching can be formulated (Birdal et al., 2021) and ambiguous candidates can be discarded during 3d reconstruction or camera pose estimation.
-3.5 Clustering as QUBO
-To solve the clustering problem using AQC, a QUBO formulation is required.
-We use a variation of the one-hot encoding approach (Date et al., 2021; Arthur & Date, 2021) that uses a matrix ${Z}\in\{0,1\}^{K\times I}$ to encode the cluster assignment of $I$ samples to $K$ clusters.
+$p({s})={\mbox{exp}[\nicefrac{{-H({s})}}{{T}}]}/{\sum_{{s}^{\prime}}\mbox{exp}[\nicefrac{{-H({s}^{\prime})}}{{T}}]},$ (4)
+where $p({s})$ describes the probability of finding the system in state ${s}$ and ${s}^{\prime}$ are all possible states.
+In this work, we use this property, to sample from the Boltzmann distribution corresponding to the energy-based model of clustering.
+4 Balanced Quantum K-Means
+In the following, we define an energy-based model (EBM) that directly connects to AQC.
+Based on this, we derive the energy function for k-means and demonstrate how an AQC can sample from the corresponding distribution.
+EBM Inference.
+We use the variable $Z\in\mathbb{Z}^{N}_{2}$, to represent our binary parameters. The posterior density of interest follows from the Bolzmann distribution as:
+$\pi(Z)\vcentcolon=p(Z|X)\propto\exp(-E(Z,X)),$ (5)
+where $X$ denotes the observations, a.k.a. the collection of all specified data and $E$ is called the potential energy:
+$E(Z,X)\vcentcolon=-(\log p(X|Z)+\log p(Z)).$ (6)
+Following [11, 12], we consider the probabilistic inference, where we will be interested in the following quantities:
+I Maximum a-posteriori (MAP):
+$\hat{Z}=\operatorname*{arg\,max}\limits_{Z\in\mathbb{Z}^{N}_{2}}\log p(Z|X)$ (7)
+where $\hat{Z}$ denotes the entirety of the sought parameters.
+II The full posterior distribution: $p(Z|X)\propto p(X,Z)$.
+Both of these problems are very challenging and cannot be directly addressed by standard methods such as gradient descent (problem I) or standard MCMC methods (problem II).
+The difficulty in these problems is mainly originated by the fact that the posterior density is non-log-concave (i.e.
+the negative log-posterior is non-convex) and any algorithm that aims at solving one of these problems should be able to operate in the particular space of binary variables.
+Nevertheless, The MAP estimate is often easier to obtain via a Quantum annealer and useful in practice [12, 7].
+On the other hand, samples from the full posterior can provide important additional information, such as uncertainty.
+Not surprisingly, the latter is a much harder task, especially considering the discrete nature of our problem.
+Unfortunately, available QAs such as D-Wave [58] are not directly capable of tackling the second problem since they sample from a modified posterior and not the true one.
+Our work fills this gap for the task of balanced K-means clustering using quantum annealing as a direct way to sample from a physical system that follows the Boltzmann distribution [68] and recomputes the posterior from the found solutions to ensure callibration.
+Clustering as QUBO.
+We now focus on obtaining a single point-estimate as a solution to problem I.
+To solve the clustering problem using AQC, a QUBO formulation of the K-means energy is required.
+We use a variation of the one-hot encoding approach [27] to formulate the QUBO of k-means, with cost terms similar to Arthur and Date [4].
+It uses a matrix ${Z}\in\{0,1\}^{K\times I}$ to encode the cluster assignment of $I$ samples to $K$ clusters.
 Each row corresponds to one of $K$ clusters and each column to one of $I$ samples.
@@
 The implementation of constrained clustering, where each cluster has a fixed size $s_{k}$ furthermore requires the row constraints $\sum_{i}{Z}_{ki}=s_{k}~{}\forall k$ on $Z$.
-With the cost $q(i,j,k)$ for assigning the pair of samples $x_{i}$ and $x_{j}$ to the same cluster $c_{k}$, the optimization problem reads
-$\hat{{Z}}=\operatorname*{arg\,min}_{Z}\sum_{k}\sum_{i}\sum_{j}Z_{ki}Z_{kj}q(i,j,k)$ (7)
-[…]
-By vectorizing $Z$ in row-major order as $\mathbf{z}=\text{vec}(Z)$, the optimization problem can be rewritten in matrix form as
-$\hat{\mathbf{z}}=\operatorname*{arg\,min}_{\mathbf{z}}\mathbf{z}^{\intercal}\,\mathbf{Q}\,\mathbf{z}\quad\text{s.t.}\quad\mathbf{G}\mathbf{z}=\mathbf{d},$ (8)
-where $\mathbf{Q}$ is a block diagonal matrix with blocks $\mathbf{Q}_{0},...,\mathbf{Q}_{K}$ and $\mathbf{G}\mathbf{z}=\mathbf{d}$ corresponds to the matrix formulation of the constraints.
-Each block $\mathbf{Q}_{k}$ of the cost matrix is a square matrix that contains the costs $q(i,j,k)$ at $\mathbf{Q}_{k,ij}$.
-While the constraints are not directly covered by the QUBO problem, they can be modeled using Lagrangian multipliers.
-To avoid a mixed discrete and continuous optimization problem, a quadratic penalty reformulation $\lambda||\mathbf{G}\mathbf{x}-\mathbf{d}||_{2}^{2}$ is chosen, leading to the minimization problem
-$\hat{\mathbf{z}}=\operatorname*{arg\,min}_{\mathbf{z}}\mathbf{z}^{\intercal}\mathbf{Q^{\prime}}\mathbf{z}+\mathbf{b^{\prime}}^{\intercal}\mathbf{z}$ (9)
-with
-$\mathbf{Q^{\prime}}=\mathbf{Q}+\lambda\mathbf{G}^{\intercal}\mathbf{G}$ (10)
-$\mathbf{b^{\prime}}=-2\lambda\mathbf{G}^{\intercal}\mathbf{b}.$ (11)
+The total energy of a solution $E(Z,X)$ can be separated into terms $e(i,j,k)$ modeling the energy of assigning the pair of samples $x_{i}$ and $x_{j}$ to the same cluster $c_{k}$.
+$E(Z,X)=\sum_{k}\sum_{i}\sum_{j}Z_{ki}Z_{kj}e(i,j,k)+E(Z)$ (8)
+where $E(Z)$ models the row and column constraints as an indicator function. By vectorizing $Z$ in row-major order as ${z}=\text{vec}(Z)$, the energy can be rewritten in matrix form as
+$E(Z,X)=E(X|Z)+E(Z)={z}^{\intercal}\,{Q}\,{z}+E(Z),$ (9)
+where ${Q}$ is a block diagonal matrix with blocks ${Q}_{0},...,{Q}_{K}$ and ${G}{z}={d}$ corresponds to the matrix formulation of the constraints.
+Each block ${Q}_{k}$ of ${Q}$ is a square matrix that contains the energy $e(i,j,k)$ at ${Q}_{k,ij}$.
+To be solved on the AQC, the potential energy needs to be formulated in a QUBO, which cannot exactly implement $E(Z)$ to model the constraints and is circumvented by using Lagrangian multipliers, leading to the MAP estimate
+$\hat{{z}}=\operatorname*{arg\,min}_{z}{z}^{\intercal}{Q^{\prime}}{z}+{b^{\prime}}^{\intercal}{z}$ (10)
+with constraints ${Q^{\prime}}={Q}+\lambda{G}^{\intercal}{G}$ and ${b^{\prime}}=-2\lambda{G}^{\intercal}{b}$.
+To avoid a mixed discrete and continuous optimization problem, a quadratic penalty reformulation $\lambda||{G}{z}-{d}||_{2}^{2}$ is chosen.
 In contrast to a linear penalty approach, where the multiplier $\lambda$ needs to be optimized, our selection only requires a sufficiently high $\lambda$, as all constraint violations result in an increased penalty term.
-With such selection, the penalty term evaluates to zero $\lambda||\mathbf{G}\mathbf{x}-\mathbf{d}||_{2}^{2}=0$ if all constraints are fulfilled, and thus, the minimizer $\hat{\mathbf{z}}$ of the modified optimization problem is a minimizer of the original optimization problem in Equation equation 8.
-Besides modeling balanced clustering, further constraints can be introduced using Lagrangian multipliers. This allows to represent a wide range of tasks as QUBO clustering problems solvable with AQC.
-4 Probabilistic Quantum Clustering
-4.1 Motivation
-Our clustering formulation employs a mixture of Gaussians to explain the observations, where the cluster assignments are latent variables.
-Each sample in a cluster $c_{k}$ is modeled as a sample drawn from a Gaussian distribution $\mathcal{N}(\mathbf{\mu}_{k},\mathbf{I})$ with mean $\mathbf{\mu}_{k}$ and identity covariance according to the k-means setting.
-Following a Bayesian approach, the best clustering solution can be found by maximizing the posterior probability over possible assignments $Z$ of points to clusters.
-$p(Z|\mathbf{X})=\frac{p(\mathbf{X}|{Z})p({Z})}{\sum_{{Z}^{\prime}}p(\mathbf{X}|{Z^{\prime}})}=\frac{p(\mathbf{X}|{Z})p({Z})}{A}.$ (12)
-While this approach provides a probabilistic estimate by jointly modeling information about the possible cluster configurations and the distribution of data points, it is often intractable to evaluate due to the partitioning function $A=\sum_{{Z}^{\prime}}p(\mathbf{X}|{Z^{\prime}})$.
-Computing $A$ requires to sum over all possible solutions $Z$, which grows exponentially with the number of samples in the clustering problem.
-To overcome this, we utilize an AQC that samples directly from a Boltzmann distribution, which we parameterize according to the probabilistic clustering problem.
-4.2 Data Model
-Determining the cost function of the optimization problem is a design choice of the algorithm.
+With such selection, the penalty term evaluates to zero if all constraints are fulfilled, and thus, the minimizer $\hat{{z}}$ of the modified optimization problem is a minimizer of the original optimization problem.
+Besides modeling balanced clustering, further constraints can be introduced using Lagrangian multipliers.
+This allows to represent a wide range of tasks as QUBO clustering problems solvable with AQC in the same way.
+4.1 Probabilistic Quantum Clustering
+The MAP estimate was obtained as the lowest energy solution measured during sampling.
+We now consider approximating the posterior distribution to quantify the confidence of these solutions matching the actual ground truth.
+Instead of relying on a sequential Markov Chain Monte Carlo (MCMC) method as done in a plethora of classical algorithms [29], we will approximate the posterior distribution by recomputing the Boltzmann distribution from the set of likely AQC solutions.
+As the likely solutions are already sampled during quantum annealing, given that the objective follows the EBM, the posterior distribution can be estimated without any large additional computational overhead.
+Note, this is unlike MCMC, which incurs significant computational load.
+Our clustering formulation employs a mixture of Gaussians to explain the observations and to define the energy function $E(X|Z)$.
+Each sample in a cluster $c_{k}$ is modeled as a sample drawn from a Gaussian distribution $\mathcal{N}({\mu}_{k},{I})$ with mean ${\mu}_{k}$ and identity covariance similar to k-means.
+As shown in Section 4, we are interested in the posterior distribution over possible assignments $Z$:
+$p(Z|{X})=\frac{p({X}|{Z})p({Z})}{\sum_{{Z}^{\prime}}p({X}|{Z^{\prime}})}=\frac{p({X}|{Z})p({Z})}{A}.$ (11)
+While this approach provides a probabilistic estimate by jointly modeling information about the possible cluster configurations and the distribution of data points, it is often intractable to evaluate due to the partition function ${\sum_{{Z}^{\prime}}p({X}|{Z^{\prime}})}$, the sum over all possible solutions $Z^{\prime}$, which incurs exponential cost in the number of samples.
+To overcome this, we utilize an AQC that samples directly from the corresponding Boltzmann distribution, which we parameterize according to the probabilistic clustering problem.
+Data Model.
+Determining the potential energy and thus, cost function of the QUBO is a design choice of the algorithm.
 For our probabilistic approach, a well-defined data distribution, that forms the basis of the cost function is required.
-While many tasks approached in quantum computer vision, such as tracking (Zaech et al., 2022) or synchronization (Benkner et al., 2021; Birdal et al., 2021), costs are based on learned metrics or heuristics, they can also be trained to reflect the properties required in our approach.
+While many tasks approached in quantum computer vision, such as tracking [77] or synchronization [6, 12], costs are based on learned metrics or heuristics, they can also be trained to reflect the properties required in our approach.
 We therefore, follow the mixture of Gaussian model, where each cluster generates samples from a normal distribution.
-With the independence of observations and clusters, given the distribution parameters, the likelihood of the joint observations for an assignment $Z$ is given as the product of the individual likelihods
-$f(\mathbf{X}|{Z})=\prod_{k=1}^{K}f(\mathbf{X}|{Z_{k}})=\prod_{k=1}^{K}\prod_{i\in Z_{k}}f(\mathbf{x}_{i}|{Z_{k}})=$ (13)
-$\prod_{k=1}^{K}\prod_{i\in Z_{k}}\frac{1}{\sqrt{2\pi^{d}}}\mbox{exp}\left[-\frac{1}{2}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})^{\intercal}\mathbf{I}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})\right]=$ (14)
-$\frac{1}{\sqrt{2\pi^{d}}}\mbox{exp}\left[-\sum_{k=1}^{K}\sum_{i\in Z_{k}}\frac{1}{2}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})^{\intercal}\mathbf{I}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})\right]$ (15)
-where the likelihood $f(\mathbf{x}_{i}|{Z_{k}})$ corresponds to a Gaussian distribution that follows $\mathcal{N}(\mathbf{\mu}_{k},\mathbf{I})$ and $d$ is the dimensionality of the space.
+With the independence of observations and clusters, given the distribution parameters, the likelihood of the joint observations for an assignment $Z$ is given as the product of the individual likelihoods
+$f({X}|{Z})=\prod_{k=1}^{K}f({X}|{Z_{k}})=\prod_{k=1}^{K}\prod_{i\in Z_{k}}f({x}_{i}|{Z_{k}})$ (12)
+where the likelihood $f({x}_{i}|{Z_{k}})$ corresponds to a Gaussian distribution that follows $\mathcal{N}({\mu}_{k},{I})$ and $d$ is the dimensionality of the space.
 This result can be used to formulate the energy-based model with the energy function
-$E(\mathbf{X}|Z)=\sum_{k=1}^{K}\sum_{i\in Z_{k}}\frac{1}{2}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})^{\intercal}\mathbf{I}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}}).$ (16)
+$E(X|Z)=\sum_{k=1}^{K}\sum_{i\in Z_{k}}\frac{1}{2}({{x}_{i}}-{{\mu}_{k}})^{\intercal}{I}({{x}_{i}}-{{\mu}_{k}}).$ (13)
 Using the energy function to rewrite the posterior distribution leads to the Boltzmann distribution from Equation 4
-$p(Z|\mathbf{X})=\frac{\mbox{exp}\left[-(E(\mathbf{X}|Z)+E(Z))\right]}{\sum_{{Z}^{\prime}}\mbox{exp}\left[-E(\mathbf{X}|Z^{\prime})\right]}.$ (17)
+$p(Z|{X})=\frac{\mbox{exp}\left[-(E({X}|Z)+E(Z))\right]}{\sum_{{Z}^{\prime}}\mbox{exp}\left[-E({X}|Z^{\prime})\right]}.$ (14)
 In this formulation, the energy of the assignment $E(Z)$ corresponds to the prior $p({Z})$, which models feasible and infeasible solutions by an indicator function.
 In the case of balanced clustering, this corresponds to allowing assignments that have each point assigned to exactly one cluster and a cluster size according to $s_{k}$.
-Searching the most likely clustering solution corresponds to finding lowest energy solution on on the AQC, as well as to the Maximum-a-Posteriori $\hat{Z}_{\text{map}}$ estimate of the assignment
-$\hat{Z}_{\text{map}}=\operatorname*{arg\,max}_{Z}p(Z|\mathbf{X})p({Z})=\operatorname*{arg\,min}_{Z}E(\mathbf{X}|Z)+E(Z).$ (18)
-As the AQC qubit system is an Ising model, it requires formulating $E(\mathbf{X}|Z)$ and $E(Z)$ quadratic in the optimization variables $Z$, enabling the joint discovery of assignments and cluster means.
-This is achieved by using the maximum likelihood estimator of the mean, resulting in a quadratic energy formulation that fits the Ising model in Equation 9
-$E_{k}(\mathbf{X}|Z)=\frac{1}{s_{k}}\sum_{i}\sum_{j}Z_{ki}Z_{kj}~{}(\mathbf{x}_{i}-\mathbf{x}_{j})^{\intercal}(\mathbf{x}_{i}-\mathbf{x}_{j}).$ (19)
-The second energy term $E(Z)$ in Equation 17, modeling the prior distribution over possible assignments cannot exactly be embedded in the Ising model.
-We therefore use the quadratic penalty method as shows in Equation equation 9 to approximate the energy of an indicator function.
-Importantly, this does not influence the energy of feasible solutions, as $\lambda||\mathbf{G}\mathbf{x}-\mathbf{d}||_{2}^{2}$ evaluates to zero for all solutions that fulfill the clustering constraints.
-4.3 Boltzmann Reparametrization
-While ideally the measurements on AQC are direct samples from the Boltzmann distribution (D-Wave Systems Inc., 2017; Dixit et al., 2021), it requires solving a range of challenges that prohibit a direct use in our scenario (Pochart et al., 2021): Mapping between the cost function and the physical system implemented on the AQC requires Hamiltonian scaling (Pochart et al., 2021).
-Estimating the scaling factor is nontrivial (Nelson et al., 2022) and prohibitive for using samples directly in many cases.
-Additionally, hardware limitations including imperfection of the processor and the spin-bath polarization effect (Pochart et al., 2021) prevent the AQC to sample the Boltzmann distribution exactly.
+When the AQC is used as an optimizer, finding the the lowest energy solution corresponds to the MAP as shown in Equation 7
+As the AQC qubit system is an Ising model, it requires formulating $E(X|Z)$ and $E(Z)$ quadratic in the optimization variables $Z$, enabling the joint discovery of assignments and cluster means.
+This is achieved by using the maximum likelihood estimator of the mean $\mu_{k}=\frac{1}{s_{k}}\sum_{i}Z_{ki}x_{i}$, resulting in a quadratic energy formulation that fits the Ising model in Equation 10
+$E_{k}({X}|Z)=\frac{1}{s_{k}}\sum_{i}\sum_{j}Z_{ki}Z_{kj}~{}({x}_{i}-{x}_{j})^{\intercal}({x}_{i}-{x}_{j}).$ (15)
+The second energy term $E(Z)$ in Equation 14, modeling the prior distribution over possible assignments cannot exactly be embedded in the Ising model and is approximated using Lagrange multipliers.
+Importantly, this does not influence the energy of feasible solutions relevant in the posterior distribution, as the penalty evaluates to zero for these.
+Posterior recomputation.
+While ideally the measurements on AQC are direct samples from the Boltzmann distribution [25, 33], it requires solving a range of challenges that prohibit a direct use in our scenario [68]: Mapping between the cost function and the physical system implemented on the AQC requires Hamiltonian scaling [68].
+Estimating the scaling factor is nontrivial [63] and prohibitive for using samples directly in many cases.
+Additionally, hardware limitations including imperfection of the processor and the spin-bath polarization effect [68] prevent the AQC to sample the Boltzmann distribution exactly.
 Finally, as the energy term $E(Z)$ can only be implemented using the penalty method, the sampling density is influenced by solutions not fulfilling the constraints.
-We compensate for these limitations by evaluating the energy of all measured feasible solutions $Z^{\prime}$ and recompute $p(Z|\mathbf{X})$ by evaluating the partitioning function over these, which only requires sampling solutions at a sufficiently high temperature.
-4.4 Maximum Pointsets
-Having the set of the most likely clustering solutions $Z$ available together with their posterior probability $P(Z|\mathbf{X})$ allows to find an assignment $Z^{*}$ of a subset of points that solves the clustering problem with increased probability $P(Z^{*}|\mathbf{X})$.
+We compensate for these limitations by evaluating the energy of all measured feasible solutions $Z^{\prime}$ and recompute $p(Z|{X})$ by evaluating the analytic partitioning function over these, which only requires sampling solutions at a sufficiently high temperature.
+Coresets.
+Having the set of the most likely clustering solutions $Z$ available together with their posterior probability $P(Z|X)$ allows to find an assignment $Z^{*}$ of a subset of points that solves the clustering problem with increased probability $P(Z^{*}|X)$.
 Such a set can be found by using Algorithm 1, which implements a greedy approach that disregards points that disagree between different clustering solutions.
+With a sufficiently well-sampled Boltzmann distribution, the resulting coreset assignment $Z^{*}$ with the corresponding probabilistic estimate $P(Z^{*}|X)$ is still calibrated.
 1:$Z^{*}\leftarrow Z_{0}$
-2:$p\leftarrow P(Z_{0}|\mathbf{X})$
+2:$p\leftarrow P(Z_{0}|{X})$
 3:$i\leftarrow 1$
@@
 8$\triangleright$ Remove points assigned to different clusters.
-9:     $p\leftarrow p+P(Z_{i}|\mathbf{X})$
+9:     $p\leftarrow p+P(Z_{i}|X)$
 10:     $i\leftarrow i+1$
@@
 12:return $Z^{*}$
-Algorithm 1 MaxsetSearch
-In the case of a sufficiently well sampled Boltzmann distribution, the resulting maxset assignment $Z^{*}$ with the corresponding probabilistic estimate $P(Z^{*}|\mathbf{X})$ is still calibrated.
-4.5 Inference Parameter Optimization
+Algorithm 1 CoresetSearch
+Lagrange multiplier optimization.
 Using the quadratic penalty method to implement the constraints requires finding suitable Lagrangian multipliers $\lambda$.
@@
 Therefore, a suitable Lagrangian multiplier lifts the cost of any constraint violation above all relevant solutions of the clustering problem, while keeping them low enough to avoid scaling the total energy of the problem up considerably.
-To estimate the multipliers, we follow an iterative procedure as also proposed in (Zaech et al., 2022).
-In an initial step, balanced k-means (Malinen & Fränti, 2014) is used to find a feasible clustering solution.
+To estimate the multipliers, we follow an iterative procedure as also proposed in [77].
+In an initial step, balanced classical k-means [56] is used to find a feasible clustering solution.
 This solution is used to estimate Lagrangian multipliers that avoid any first-order violation of the constraints.
-In subsequent iterative optimization steps, the problem is solved in simulation, to find multipliers that result in a well conditioned problem.
-Such optimization procedure is crucial due to the low fidelity of the current generation of AQCs, which requires careful engineering of the problem energy.
+In subsequent iterative optimization steps, the problem is solved in simulation, to find multipliers that result in a well-conditioned problem.
+Such optimization procedure is crucial due to the low fidelity of the current generation of QAs, which requires careful engineering of the problem energy.
 Therefore, we expect this procedure to become of reduced importance with the progress of quantum computing.
 5 Experiments and Results
-We perform experiments on synthetic data as well as real data to verify the efficacy of our method in finding the set of high-probability solutions and in estimating calibrated confidence scores.
-The experimental scenarios are solved with QA, Simulated Annealing (SIM), and exact exhaustive search using the presented energy formulation and with k-means as a baseline method, which all optimize for the same cost objective.
+We perform experiments on synthetic as well as real data to verify the efficacy of our method in finding the set of high-probability solutions and in estimating calibrated confidence scores.
+The experimental scenarios are solved with QA, Simulated Annealing (SIM), and exact exhaustive search using the presented energy formulation and with k-means as a baseline method.
 This further allows us to understand the limitations and required work when deploying the approach to real quantum computers.
-5.1 Solver Methods
+Figure 2: Evaluation of the calibration and distribution for QA, SIM and exhaustive search in tasks with 2 to 4 clusters, each with 5 points.
+All results are generated with 1,000 problems in each scenario and 5,000 measurements for each clustering problem.
+Implementation details.
 […]
-Due to the strong compute-time limitations on an AQC, all Lagrangian optimization steps are performed with SIM, before measuring the final results on the AQC.
 Simulated annealing (SIM) provided by D-Wave is used for larger scale comparisons.
@@
 By iterating all feasible solutions, the lowest energy solution is guaranteed to be found and the partitioning function is computed exactly.
-K-Means clustering with a balanced cluster constraint (Malinen & Fränti, 2014) forms the baseline for our approach.
+K-Means clustering with a balanced cluster constraint [56] forms the baseline for our approach.
 We run the algorithm until convergence for a maximum of 1000 iterations.
 While this solution does not provide a probabilistic estimate, it is useful to assess the relative clustering performance.
-5.2 Dataset and Metrics
-Data for the quantitative evaluation of our method is synthetically generated.
-For each clustering problem a total of $I$ points are sampled from a separate normal distribution for each of $K$ clusters.
-The cluster centers are randomly drawn, such that the distance between each pair of clusters lies within a predefined range $[d_{\text{min}},d_{\text{max}}]$.
-For each experiment a total of $L$ clustering tasks is generated.
-This allows us to evaluate the calibration metrics over a large value range.
-For all experiments that directly compare methods, identical clustering tasks are used.
-Further qualitative examples are provided for the IRIS dataset (Fisher, 1936), which contains 50 samples of 4 features in 3 classes.
-We randomly subsample the points and dimensions to generate the parameters required for our experiments.
-Clustering metrics are computed using the available ground-truth clusters.
-We evaluate 4 standard metrics: The accuracy, which measures the ratio of clustering solutions that are identical to the ground-truth.
-Completeness (Rosenberg & Hirschberg, 2007) measures the ratio of points from a single cluster being grouped together.
-The adjusted Rand score (Hubert & Arabie, 1985) compares all pairs of points in the ground truth and prediction and the Fowlkes-Mallows index (Fowlkes & Mallows, 1983) combines precision and recall into a single score.
 Solver
@@
 Table 1: Synthetic data results for our approach (QA, SIM, exhaustive) and k-means. All numbers in % with standard error of the mean.
-5.3 Calibration Performance
+Data for the quantitative evaluation of our method is synthetically generated.
+For each clustering problem a total of $I$ points are sampled from a separate normal distribution for each of $K$ clusters.
+The centroids are randomly drawn, such that the distance between each pair of clusters lies within a predefined range $[d_{\text{min}},d_{\text{max}}]$.
+For each experiment a total of $L$ clustering tasks is generated.
+This allows us to evaluate the calibration metrics over a large value range.
+For all experiments that directly compare methods, identical clustering tasks are used.
+Further results are provided for the IRIS dataset [38] and a dataset of images containing ambiguous objects that are to be identified.
+The IRIS contains 50 samples of 4 features in 3 classes.
+We randomly subsample the points and dimensions to generate the parameters required for our experiments.
+Clustering metrics are computed using the available ground-truth clusters.
+We evaluate 4 standard metrics: The accuracy, which measures the ratio of clustering solutions that are identical to the ground-truth.
+Completeness [69] measures the ratio of points from a single cluster being grouped together.
+The adjusted Rand score [44] compares all pairs of points in the ground truth and prediction and the Fowlkes-Mallows index [40] combines precision and recall into a single score.
+5.1 Results
+Calibration performance.
 We evaluate the calibration of our method on a synthetic clustering scenario with 3 clusters and 15 points using QA, SIM and exhaustive search on 1000 tasks.
-First all clustering solutions $Z$ are accumulated in bins according to their estimated posterior probability $P(Z|\mathbf{X})$.
+First all clustering solutions $Z$ are accumulated in bins according to their estimated posterior probability $P(Z|X)$.
 This process also includes all sampled non-optimal but feasible solutions.
-The resulting histogram of solutions is shown in Figure 1(a).
-After accumulation, the ratio of correct solutions in each bin is evaluated in Figures 1(b),1(c), and 1(d), which for a calibrated method should be close to the mean predicted probability represented by the diagonal.
-We find our approach to generate well-calibrated probabilities in both simulation and when using QA.
-(a) Histogram of predicted probabilities.
-(b) 2 clusters, 10 points.
-(c) 3 clusters, 15 points.
-(d) 4 clusters, 20 points.
-Figure 2: Evaluation of the calibration for QA, SIM and exhaustive search in a clustering scenario with 15 points and 3 clusters.
-Evaluation of the calibration for quantum annealing, simulated annealing and exhaustive search in clustering scenarios with 2, 3, and 4 clusters and 5 points in each cluster.
-All results are generated with 1,000 problems in each scenario and 5,000 measurements for each clustering problem.
-5.4 Clustering Performance
+Figure 2 shows the resulting histogram of solutions and the ratio of correct solutions in each bin after accumulation.
+It thus evaluates calibration, where the ideal case has mean predicted probabilities on the diagonal.
+We find our approach to generate well-calibrated probabilities in both simulation and when using QA for the two smaller experiments for up to 15 points.
+Due to the problem QUBO size, only SIM is able to find the correct solutions on the largest example.
+Clustering performance.
 We evaluate the performance of our clustering formulation using synthetic data in Table 1.
 The upper rows show results for 3 clusters with an increasing number of points and 1000 tasks/setting for SIM and k-means.
-Due to the problem size, results cannot be found using QA and exhaustive search.
-Comparing results shows that most clustering metrics are better when using our formulation with SIM, however, the difference decreases with increasing problem size.
+Due to the problem size, it cannot be solved using QA and exhaustive search.
+Our formulation with SIM outperforms k-means on the small tasks, however, the difference vanishes with increasing problem size.
 This can be attributed to the globally optimal solution our approach is optimizing for.
-While in an ideal the global solution is at least as good as the k-means solution, an increasing problem size also increases the complexity of the corresponding QUBO.
-This makes it harder to solve exactly and the noisy approach using annealing is not capable of solving the problem correctly.
-For the largest clustering problem with 3 clusters and 45 points, k-means provides the best accuracy and thus, a larger number of correct solutions compared to SIM.
-However, the metrics indicating clustering quality are higher for SIM.
+Ideally, the global solution is at least as good as the k-means solution, however, an increasing problem size also increases the complexity of the QUBO, which explains the dropping performance on large problems.
+For the largest clustering problem with 3 clusters and 45 points, k-means provides the best accuracy with more correct solutions compared to SIM.
+However, the clustering quality metrics are higher for SIM.
 This indicates that our formulation is able to find a better solution in cases where the problem is not solved correctly by SIM and k-means.
@@
 The quality of predicted posterior probabilities can be further evaluated by linking them to the clustering metrics and sparsifying the set of tasks.
-Starting with metrics over the whole set of clustering tasks, we remove tasks according to increasing predicted probability and evaluate the metric over the remaining set.
-In an ideal predictor, this removes the lowest performing tasks first.
-In Figures LABEL:fig:sparsification_method_comp and LABEL:fig:sparsification_metric_comp this is done for the adjusted Rand index with different solvers and for QA with different metrics respectively.
-The solid lines show the sparsification-plots using the predicted probabilities and the dashed lines show the same plots for an oracle method that generates the best possible ordering based on the metrics themselves.
+Starting with metrics over the whole set of clustering tasks, we remove tasks according to increasing probability and evaluate the metric over the remaining set.
+For an ideal predictor, this removes the lowest-performing tasks first.
+In Figure LABEL:fig:sparsification_method_comp this is done for the adjusted Rand index with different solvers and in Figure LABEL:fig:sparsification_metric_comp QA with different metrics.
+The solid lines show the sparsification plots using the predicted probabilities and the dashed lines show the same plots for an oracle method that generates the best possible ordering based on the metrics.
 In Figure LABEL:fig:sparsification_method_comp it becomes apparent that QA, SIM and exhaustive search perform close to each other over most of the value range.
 Nevertheless, for a high sparsification with more than 80% of the tasks removed, QA shows a drop in performance compared to the other methods.
-This is caused by tasks where only a single, but incorrect solution is found, which gets assigned a posterior probability of $P(Z|\mathbf{X})=1.0$.
-In such cases the Boltzmann distribution has not been sampled sufficiently well, either because of too few measurements or because of a low effective sampling temperature.
+This is caused by tasks where only a single, but incorrect solution is found, which gets assigned a posterior probability of $P(Z|X)=1.0$.
+In such cases, the Boltzmann distribution has not been sampled sufficiently well, either because of too few measurements or because of a low effective sampling temperature.
 As SIM does not show this behavior the source can likely be traced back to current limitation of the quantum computer.
-5.5 Maximum Pointsets
-Qualitative examples for the maximum pointsets generated with Algorithm 1 are depicted in Figure 4.
+Coresets.
+Qualitative examples for the coresets generated with Algorithm 1 are depicted in Figure 4.
 The shape of each point represents the ground truth class and the color the assigned cluster.
-Starting with the most likely solution having a predicted probability of $p(Y|X)=0.61$ on the left, each plot shows one additional step of the algorithm, which successively removes points from the solution, indicated by plotting them in Grey.
-The illustration demonstrates that the MaxsetSearch algorithm is able to generate well-separated clusters from the probabilistic predictions.
+Starting with the most likely solution having a predicted probability of $p(Z|X)=0.61$ on the left, each plot shows one additional step of the algorithm, which successively removes points from the solution, indicated by plotting them in Grey.
+The illustration demonstrates that the CoresetSearch algorithm is able to generate well-separated clusters from the probabilistic predictions.
 (a) $p=0.61$
@@
 (d) $p=0.90$
-Figure 4: Visualization of max pointsets for synthetic data with the probability for each of the determined pointsets.
-5.6 IRIS Dataset
-While our method assumes an identity covariance in the data, it can be applied to other distributions, which we evaluate on the widely used IRIS dataset.
-Clustering metrics for experiments using 3 clusters with 5 points each are provided in Table 6 and show that our formulation using QA and SIM is competitive with k-means.
+Figure 4: Visualization of coresets for synthetic data with the probability for each of the determined pointsets.
+{floatrow}
+[…]
+[…]
+Figure 6: Performance on IRIS subsets (3 clusters/15 points).
+Figure 5: Qualitative results on the IRIS dataset.
+[…]
+[…]
+Real-World Datasets
+While our method assumes an identity covariance in the data, it can be applied to other distributions, which we evaluate on the widely used IRIS dataset as well as a set of self-collected images.
+Clustering metrics for experiments on IRIS using 3 clusters with 5 points each are provided in Table 5(b) and show that our formulation using QA and SIM is competitive with k-means.
 Figures LABEL:fig:qual_iris_qa and LABEL:fig:qual_iris_sim show differently sized qualitative examples from the dataset solved using our formulation with QA and SIM respectively.
-[…]
-Figure 5: Qualitative results on the IRIS dataset.
-Figure 6: Clustering metrics generated on IRIS subsets using 3 clusters with 15 points in total.
-Though the value of the predicted probability cannot be interpreted easily with the mismatch between the assumed and the actual data-generating process, it can be used to find maximum pointsets.
-The results from Algorithm 1 using results from SIM are provided in Figure 7 and show that even for a distribution mismatch the maximum pointsets can provide meaningful results for removing ambiguous samples.
-Figure 7: Max pointsets generated using simulated annealing on the IRIS dataset with 60 points.
+On the full IRIS dataset, SIM and k-means achieve identical results with a Completeness of 77.7%, Adjusted Rand index of 78.6% and further results provided in the supplementary.
+The results from Algorithm 1 using results from SIM are provided in Figures 7 and 8 for IRIS and our dataset respectively.
+They show that even for a distribution mismatch the coresets can provide meaningful results for removing ambiguous samples.
+To demonstrate that meaningful ambiguous samples can be identified in a high-dimensional space present in image data, we collected publicly available images of cars, boats and cars towing boats.
+Image features were extracted using VGG [72] and subsequently clustered using our approach.
+Similar to IRIS, the distribution does not match our assumptions, however, we are still able to identify ambiguous images that contain cars towing a boat by using Algorithm 1.
+The first coreset is depicted in Figure 8 where differently colored frames indicate the cluster assignment and Grey frames mark ambiguous samples.
+Clustering is performed on the 1000-dimensional feature and images are plotted at their 2D projection generated using UMAP [59].
+Our approach correctly identifies the images that contain a car towing a boat and thus, cannot be uniquely assigned to either cluster.
+Figure 7: Coresets generated using SIM on the IRIS dataset with 60 points.
+Figure 8: First coreset using SIM on our collected dataset.
 6 Conclusion
@@
 By using all valid measurements, calibrated confidence scores are computed at little cost and solutions are competitive to an iterative balanced k-means approach.
-We evaluated our method on synthetic as well as real data using simulation, exhaustive search as well as the D-Wave Advantage 2 prototype AQC to explore the potential of quantum computing in machine learning and computer vision.
+We evaluated our method on synthetic as well as real data using simulation, exhaustive search as well as the D-Wave Advantage 2 prototype QA to explore the potential of quantum computing in machine learning and computer vision.
 While the approach is still limited in the problem size, quantum computing enables a fundamentally different approach to clustering that can provide additional information that is costly to compute otherwise.
 Nevertheless, even with the current progress and potential to scale to real-world problems, more work is required to adapt existing problem formulations, such that the full capability of quantum computing can be effectively used.
-Acknowledgments
+Acknowledgement:
 This work was funded by Toyota Motor Europe via the research project TRACE Zürich.